Hi please help me with this problem: If line segment BC bisects angle ABD and angle ABD=106 degrees, find x and y.
angle ABC=(5x-y)
angle CBD=(4x+2y+19)

Answer by stanbon(75887)

If line segment BC bisects angle ABD and angle ABD=106 degrees, find x and y.
-----

angle ABC=(5x-y) = 53 degree
angle CBD=(4x+2y+19) = 53 degree
-------------------
5x - y = 53
4x + 2y = 34
------
Modify for elimination:
5x - y = 53
2x + y = 17
------
7x = 70
x = 10
---
Solve for "y"::
5x - y = 53
y = 5x - 53
y = 50 - 53
y = -3
------------------
Cheers,
Stan H.
-------------------